OpenStudy (albert0898):

Tanya has ten bills in her wallet. She has a total of $40. If she has one more $5 bill than $10 bills, and two more $1 bills than $5 bills, how many of each does she have? (There are two ways of working this problem. See if you can do it both ways.) EXPLAIN IT and SHOW STEPS PLEASE.

OpenStudy (anonymous):

Straight algebraic (x,x+1,x+3, etc) or Logical deduction (Must be less than 4 tens etc)

Parth (parthkohli):

\(\Large \color{purple}{5(x + 1) + 10x + 1(x + 3) }\) \(\Large \color{purple}{ 5x + 5 + 10x + x + 3 }\) \(\Large \color{purple}{ 16x + 8 = 40 }\) \(\Large \color{purple}{ 8(2x + 1)=40 }\) \(\Large \color{purple}{ 2x + 1 = 5 }\) \(\Large \color{purple}{ 2x = 4 }\) \(\Large \color{purple}{ x = 2 }\) Now, there are 3 $5 notes, 2 $10 notes, 5 $1 notes

Parth (parthkohli):

Can you understand how this went? @Albert0898

OpenStudy (albert0898):

Not really.

Parth (parthkohli):

Hmm, see, there are 10 notes, that doesn't really matter. See, so she has one more $5 note than the $10 notes. To calculate money, we must multiply the money by the number of notes. Simple. 5(x + 1) + 10(x) + 1(x + 3) = 40 Solve it now.

OpenStudy (albert0898):

5x + 5 + 10x + 1x + 3 = 40 16x + 8 = 40 16x = 32 x = 2 Can you just explain to me this sentence... "she has one more $5 bill than $10 bills, and two more $1 bills than $5 bills"?

OpenStudy (albert0898):

@ParthKohli

Parth (parthkohli):

Okay, so like if you have six $5 bills and five $10 bills, then you have one more $5 bill than you have $10 bills.

Parth (parthkohli):

See the answer, we have one more 5 dollar note than the number of 10 dollar notes.

OpenStudy (anonymous):

BILLS POBLEMS Let X = 10 dollar bills then , the 5 dollar bills = X +1 and the 1 dollar bills = X + 3 Now, you have to multilpy each expression by its corresponding monetary value, Equation : 10 X + 5 (X +1) + X + 3 = 40 16 X = 32 X = 2 So , 10 dollar bills = 2 5 dollar bills = 2+ 1 = 3 1 dollar bills = 2 + 3 = 5 Verification : ($10)2 + ($5)3 + ($1)5 = 40 $20 + $15 + $5 = $40